OpenStudy (anonymous):

Finish these three solutions to this equation using the given value of x. y = 2x -3 (1,____), (2, ____), (3, ____)

jimthompson5910 (jim_thompson5910):

if x = 1, then... y = 2x -3 y = 2*1 -3 ... replace x with 1 (since x = 1) y = 2 - 3 y = -1 So if x = 1, then y = -1 The first ordered pair is (1, -1)

jimthompson5910 (jim_thompson5910):

The other ordered pairs are found in the same way